### 30 Jan 2024 v2

# Data Distribution Bottlenecks in Grounding Language Models to Knowledge Bases

Abstract
Language models (LMs) have already demonstrated remarkable abilities in understanding and generating both natural language and formal language. Despite these advances, their integration with real-world environments such as large-scale knowledge bases (KBs) remains an underdeveloped area, affecting applications such as semantic parsing and indulging in "hallucinated" information. This paper is an experimental investigation aimed at uncovering the robustness challenges that LMs encounter when tasked with knowledge base question-answering (KBQA). The investigation covers scenarios with inconsistent data distribution between training and inference, such as generalization to unseen domains, adaptation to various language variations, and transferability across different datasets. Our comprehensive experiments reveal that even when employed with our proposed data augmentation techniques, advanced small and large language models exhibit poor performance in various dimensions. While the LM is a promising technology, the robustness of the current form in dealing with complex environments is fragile and of limited practicality because of the data distribution issue. This calls for future research on data collection and LM learning paradigms [^1] Code and data will be public upon acceptance..
Data Distribution Bottlenecks in Grounding Language Models to Knowledge Bases
Yiheng Shuthanks: * Work performed during internship at Microsoft Research Asia. The Ohio State University shu.251@osu.edu                        Zhiwei Yu Microsoft zhiwyu@microsoft.com
1 Introduction
Language models (LMs), such as BERT Devlin et al. (2019), T5 Raffel et al. (2020), and the GPT series Ouyang et al. (2022); OpenAI (2023), have demonstrated impressive capabilities in understanding and generating natural or formal language, highlighting the potential for artificial general intelligence (AGI). However, several obstacles must be overcome to achieve this goal. For example, to mitigate the issue of "hallucinated" information and expand the range of LM applications, researchers have focused on grounding LMs to real-world environments (e.g., Web, database, knowledge base) Nakano et al. (2021); Menick et al. (2022), which enables real-time fact-checking, data validation and thereby improve the reliability of model responses. Knowledge base (KB) is one underdeveloped environment among them. The task of Knowledge Base Question Answering (KBQA) aims to parse natural language queries using KBs, such as Freebase Bollacker et al. (2008), and Wikidata Vrandecic and Krötzsch (2014). Now, numerous LM-driven models Das et al. (2021); Hu et al. (2022) continue to achieve higher scores on KBQA benchmarks, but grounding LMs to KBs has not been fully examined for robustness and reliability. A few critical gaps in existing works prompt the necessity for this work. First, while the evaluation of LMs typically occurs in natural language tasks, the challenge escalates when grounding these models to real-world environments like KBs Liu et al. (2023), where the data contains structured data instead of all unstructured natural language. Second, the metrics of the KBQA benchmark are often shallow and the robustness of the model is not adequately evaluated. Finally, recent surveys on KBQA Lan et al. (2022); Gu et al. (2022b) have largely omitted the strides made in the development and application of LMs, especially large language models (LLMs). As a result, there remains a clear gap in understanding the robustness challenges and solutions specific to grounding LMs to KBs.
Figure 1: Inconsistent KBQA data distribution between training and inference.
For deep learning models, robustness is closely related to data distribution Hendrycks et al. (2020). The efficacy of language models (LMs) is not just a product of their architecture, but also a reflection of the data on which they are trained. In simpler tasks and well-defined domains, large-scale corpora have been collected and used for effective training Touvron et al. (2023). However, real-world environments are rarely so accommodating, e.g., large KBs contain complex structures and schema items, and building a representative corpus is more challenging. Inconsistency in data distribution during training and inference, as shown in Figure 1, may negatively impact the performance and robustness of LMs. Given this backdrop, it becomes essential to approach the task of grounding LMs from a multifaceted standpoint. Specifically, we believe that a detailed understanding of this grounding task requires exploring environmental, linguistic, and model learning aspects.
Through these explorations, this paper aims to provide a more holistic understanding of the challenges and opportunities that arise in the LM grounding and KBQA benchmarking processes. We review existing works and identify several challenges, including: 1) generalization to unseen domains at the schema level Gu et al. (2021), 2) adaptation to paraphrases featuring diverse language variations Su et al. (2016), 3) transferability across datasets with novel schema items, query patterns, and linguistic styles Cao et al. (2022b), and 4) few-shot in-context learning capabilities of grounding LLMs Li et al. (2023). These aspects allow us to dissect and critique the robustness and applicability of LMs when operating in complex real-world scenarios, thereby fulfilling our objective of effective interaction between LMs and environments.
To substantiate the presence of these challenges and obtain additional empirical insights, we undertake a comprehensive experimental investigation for the above aspects. To mitigate the negative effects of inconsistent data distributions in our experiments, we propose a data augmentation method for any LM and a retrieval augmentation method for any LLM. Our findings reveal that even when employed with such techniques (the highest EM score is achieved on the GrailQA benchmark), advanced small and large LMs still fall short of effectively tackling the majority of these challenges. A striking example is the large difference between the best practice without WebQSP Yih et al. (2016) fine-tuning (F1 43.0%) compared to the fine-tuned state-of-the-art (F1 79.6%) (§6), suggesting the weak robustness of LM-driven KBQA models on an unseen dataset. Such observations highlight an urgent need for future research in data collection methodologies and LM learning paradigms. Meanwhile, we expect the evaluation approaches in this paper to provide a reference for future benchmark construction, developing metrics that take robustness into account (§3).
2 Preliminary
In this paper, the knowledge base (KB) refers to an RDF [^2] Resource Description Framework by W3C standard. graph, consisting of triples $(s,r,o)$, where $s$ is a subject, $r$ is a relation and $o$ is an object. Logical form refers to a formal query that can be executed on a KB like SPARQL and S-expression Gu et al. (2021). Schema refers to the rdfs:Class (class) and rdf:Property (relation) in the RDF Schema [^3] https://www.w3.org/TR/rdf12-schema/.
Benchmark KB Size LF Generalization Para.
WebQuestions Berant et al. (2013) Freebase 5,810 N/A i.i.d.
SimpleQuestions Bordes et al. (2015) Freebase 108,442 N/A i.i.d.
WebQuestionsSP Yih et al. (2016) Freebase 4,737 SPARQL i.i.d.
GraphQuestions Su et al. (2016) Freebase 5,166 Graph query comp.+zero
LC-QuAD Trivedi et al. (2017) DBpedia 5,000 SPARQL i.i.d.
CWQ Talmor and Berant (2018) Freebase 34,689 SPARQL i.i.d.
LC-QuAD 2.0 Dubey et al. (2019) Wikidata 30,000 SPARQL i.i.d.
SQB Wu et al. (2019) Freebase 108,443 N/A i.i.d.+zero
CFQ Keysers et al. (2020) Freebase 239,357 SPARQL comp
GrailQA Gu et al. (2021) Freebase 64,331 S-expression i.i.d.+comp.+zero
KQA Pro Cao et al. (2022a) Wikidata 117,970 KoPL i.i.d.
QALD series Perevalov et al. (2022) DBpedia 558 SPARQL comp.
Table 1: Selected KBQA benchmarks. LF: logical forms. Generalization settings follow Gu et al. (2021). i.i.d. denotes that the schema distribution in the test set is the same as the training set. comp. and zero denote compositional and zero-shot generalization, respectively. Para. denotes paraphrases, i.e., questions containing the same semantics (machine-generated paraphrases are not included).
3 Robustness Challenge
In this paper, the robustness of a KBQA model refers to its ability to adapt to various natural language inputs and maintain consistent performance when data distribution shifts. Due to the data discrepancy between the training corpus of LMs and KB environments, LMs face challenges from environmental, linguistic, and model learning aspects.
3.1 Environmental Aspect
One of the principal challenges from an environmental perspective is schema-level generalization. The RDF Schema offers a data-modeling vocabulary, which is essential for querying a KB. As shown in Table 1, most KBQA benchmarks operate under the assumption that the schema distribution is identically and independently distributed (i.i.d.) between training and testing scenarios. This means that the schema items encountered during testing are already familiar to the model from the training data. However, this assumption often falls short in large-scale KBs, which contains thousands of schema items. Currently, only a limited number of benchmarks aim to evaluate the ability to generalize at the schema level, especially when it comes to handling non-i.i.d. schema items. The SimpleQuestions-Balance dataset Wu et al. (2019) reconfigures the i.i.d. SimpleQuestions dataset Petrochuk and Zettlemoyer (2018) to ensure that half of the questions in the testing (or validation) set are mapped to KB relations not present in the training data. GrailQA Gu et al. (2021) is structured to evaluate three levels of schema-level generalization: i.i.d. (25%), compositional generalization (25%), and zero-shot generalization (50%). GraphQuestions Su et al. (2016) presents an even more rigorous test as its testing set features schema items seldom encountered in the training data. Given the impracticality of maintaining the i.i.d. assumption in real-world applications, these datasets offer a more realistic portrayal of the generalization challenges that QA systems face in practical settings. Despite recent advances Shu et al. (2022); Gu et al. (2022a), challenges in compositional and zero-shot generalization are far from solved.
3.2 Linguistic Aspect
Natural language is variable in form, making question understanding challenging for KBQA models. One common way this variety shows up is through paraphrasing. In this paper, a paraphrase set denotes different ways to express the same logical form, as illustrated in Table 14. To gauge how well a KBQA model can adapt to these different forms of natural language, a straightforward approach is to test if the model can accurately answer paraphrased questions that it has already answered correctly before. Unfortunately, as shown in Table 1, many KBQA benchmarks do not account for paraphrasing with only one utterance for each logical form. Exceptionally, some datasets Su et al. (2016); Dubey et al. (2019); Gu et al. (2021) are based on automatically generated logical forms and include multiple natural language expressions for the same logical form (template). These data characteristics highlight the difficulties in adapting to paraphrased questions.
3.3 Integrated Aspect
Evaluating KBQA benchmarks often hinges on a single dataset, thereby complicating the task of ascertaining the model performance consistency across novel scenarios. This form of robustness, termed as cross-dataset transfer in this paper, combines both the environmental and linguistic aspects discussed earlier and is more difficult to achieve. This is because construction methods vary across datasets, as do schema distributions and natural language expressions. Specifically, KBQA dataset construction generally falls into two distinct categories: 1) Graph Search and Crowdsourcing: in this approach, logical forms or triples are initially extracted from a KB, where structures or operators of logical form are usually finite. Subsequently, these are converted into natural language utterances through crowdsourcing techniques Bordes et al. (2015); Trivedi et al. (2017). 2) Human Curation and Parsing: logical forms are labeled directly from human-provided utterances Berant et al. (2013); Perevalov et al. (2022). Existing works Gu et al. (2021); Cao et al. (2022b) suggest that models pre-trained on large-scale datasets can adapt reasonably well to other target datasets, such as WebQuestionsSP Yih et al. (2016). However, the necessity for fine-tuning these pre-trained models on the intended target dataset remains imperative for achieving optimal performance. That is, despite the advantages offered by pre-training on expansive KBQA datasets, models still encounter challenges in transferring directly to previously unseen target datasets while sustaining high performance.
3.4 Learning Aspect
Aside from considering environmental and linguistic factors, it is crucial to focus on the model learning process. Recently, LLMs like GPT series OpenAI (2023) have demonstrated exceptional capabilities across a variety of tasks, outperforming smaller yet potent LMs such as BERT Devlin et al. (2019) and T5 Raffel et al. (2020). Despite these advancements, these LLMs face substantial challenges when interacting with environments. One notable issue is their predominant reliance on an in-context learning paradigm as opposed to fine-tuning, as a trade-off between computational cost and model efficiency. In comparison to fine-tuning, in-context learning offers the advantage of reduced training costs but at the expense of the lacking perception of the environment. Inconsistent data distribution between natural language pre-training and reasoning over structured knowledge contexts leads to poor performance. For instance, a discernible performance gap exists between KBQA models that employ in-context learning with Codex Chen et al. (2021a) and those built on fine-tuned LMs Gu et al. (2022a); Li et al. (2023). Therefore, it is crucial to consider the limitations of the commonly used in-context learning paradigm and improve the grounding approaches.
4 Approach
The goal of our proposed approaches is to reduce the negative effects of inconsistent data distributions and to maximize the potential of LMs in our experimental investigations. We introduce both data augmentation and retrieval augmentation techniques.
4.1 Data Augmentation for LMs
Off-the-shelf datasets of limited size may make LM easily overfitted and not adaptable to large KBs. To address the problem that many domains in the KB are often not collected as training data, we propose a data augmentation method named Graph seArch and questIon generatioN (GAIN). GAIN applies to KBQA corresponding to logical forms or triples, and scales data volume and distribution through four steps: 1) Graph search: Sampling logical forms or triples from arbitrary domains in the KB, without being restricted to any particular KBQA dataset. 2) Training question generator on existing KBQA datasets, i.e., learning to convert logical forms or triples into natural language questions. 3) Verbalization: Using the question generator from step 2 to verbalize sampled logical forms or triples from step 1, thus creating synthetic questions. 4) Training data expansion: Before fine-tuning any neural models on KBQA datasets, GAIN-synthetic data can be used to train these models or to expand the corpus of in-context samples for LLMs. That is, as a data augmentation method, GAIN is not a KBQA model, but it is used to augment a base KBQA model.
4.2 Retrieval Augmentation for LLMs
As the trade-off between cost and effectiveness, we experiment with the prevalent in-context learning paradigm but attempt to improve the quality of in-context samples. We use advanced retrieval methods based on smaller LMs as plug-ins to augment the LLM, similar to the SuperICL approach Xu et al. (2023). Specifically, our steps to generate an LLM prompt for each question include the following. 1) Given an input question, we retrieve $k$ questions ($k$-shot) with BM25 Robertson et al. (2009) from the corpus (the combination of KBQA training set and the GAIN-synthetic dataset). 2) The role of retrieval augmentation for KB environments has been shown by fine-tuned LMs Shu et al. (2022). To assist with grounding LLM, we retrieve KB contexts with off-the-shelf retrievers for $k$ samples and the input question [^4] The prompt example is demonstrated in Appendix A..
5 Experiments
5.1 Setup
Datasets
We use GrailQA Gu et al. (2021) and GraphQuestions Su et al. (2016) for schema-level generalization and paraphrase settings. We also study generalization on SimpleQuestions-Balance (SQB) Wu et al. (2019) with unseen relations. WebQuestionsSP (WebQSP) Yih et al. (2016) is employed for cross-dataset transfer experiments as it is based on practical human-curated search logs. All experiments use S-expression as the logical form due to its clear and concise structure. Entity linking results are taken from TIARA Shu et al. (2022) for GrailQA and WebQSP, and ArcaneQA Gu and Su (2022) for GraphQuestions.
Model
We report the performance of comparison models from their papers. For the relation linking task on SQB, we use BERT Devlin et al. (2019) as the base model for GAIN. For KBQA tasks, we use the open-source advanced model TIARA Shu et al. (2022) as the base model for GAIN, due to its strong performance on zero-shot schema items [^5] Pangu Gu et al. (2022a) also uses entity linking results from TIARA.. TIARA is composed of multi-grained retrievers and a generator, with the retrievers providing KB contexts [^6] Entities, exemplary logical forms (ELF) and schema items are retrieved. for the generator. The term “TIARA+GAIN” represents a model (both the retrievers and the generator) that is first tuned using GAIN synthetic data and subsequently fine-tuned on a target dataset. For LLM evaluation, we use the latest gpt-3.5-turbo-0613777https://platform.openai.com/docs/models model, and the few-shot contexts are retrieved from the combination of GrailQA training set and synthetic dataset using the TIARA+GAIN retrievers.
Metrics
Following previous works, we use Exact Match (EM), F1, and Hits@1 to measure the performance of KBQA models. To evaluate adaptability to paraphrases (§3.2), we calculate the standard deviation (std) of EM/F1 for questions of each logical form template. As shown in Equation 1, suppose there are $n$ sets of paraphrases in the dataset, each set of paraphrases corresponds to a logical form template with $m$ natural language expressions, and the F1 score obtained by the KBQA model on the $j$-th question of the $i$-th set of paraphrases is $F1_{i,j}$. The metric $Std_{F1}$ first calculates the standard deviation of the F1 scores obtained by the model on the $m$ questions for each set of paraphrases and then calculates the average of the $n$ standard deviations. This metric is used to measure the robustness of the model to different representations of the same semantics, i.e., whether it can cope with diverse natural language expressions. A lower standard deviation indicates that the model is more adaptive to different expressions. $Std_{EM}$ is calculated in the same way.
$Std_{F1}=\frac{1}{n}\sum_{i=1}^{n}\sqrt{\left(\frac{\sum_{j=1}^{m}(F1_{i,j}-\bar{F1_{i}})^{2}}{m}\right)}$ (1)
#question #one-hop #two-hop #domain
127,329 78,668 48,661 759
#none #count #comparatives #superlatives
115,221 7,115 1,874 3,119
#class #relation #entity
5,078 12,942 46,645
Table 2: Statistics for the synthetic dataset of logical forms. none denotes no function.
#question #relation #subject #domain
162,557 7,349 108,804 673
Table 3: Statistics for the synthetic dataset of triples. Subject denotes subject entities.
Overall I.I.D. Compositional Zero-shot
Model on GrailQA Test Set EM F1 EM F1 EM F1 EM F1
Fine-tuned Models
BERT + Ranking (Gu et al., 2021) 50.6 58.0 59.9 67.0 45.5 53.9 48.6 55.7
RnG-KBQA (Ye et al., 2022) 68.8 74.4 86.2 89.0 63.8 71.2 63.0 69.2
TIARA (T5-base) (Shu et al., 2022) 73.0 78.5 87.8 90.6 69.2 76.5 68.0 73.9
DecAF (FiD-3B) (Yu et al., 2022) 68.4 78.8 84.8 89.9 73.4 81.8 58.6 72.3
Pangu (BERT-base) (Gu et al., 2022a) 73.7 79.9 82.6 87.1 74.9 81.2 69.1 76.1
Pangu (T5-large) (Gu et al., 2022a) 74.8 81.4 82.5 87.3 75.2 82.2 71.0 78.4
Pangu (T5-3B) (Gu et al., 2022a) 75.4 81.7 84.4 88.8 74.6 81.5 71.6 78.5
Codex-driven Models
B-BINDER (6)-R  (Li et al., 2023) 53.2 58.5 72.5 77.4 51.8 58.3 45.0 49.9
Pangu (Codex) (Gu et al., 2022a) 56.4 65.0 67.5 73.7 58.2 64.9 50.7 61.1
GAIN-augmented Models
TIARA + GAIN (T5-base) 75.1 80.6 88.3 91.0 73.0 79.6 69.9 76.4
TIARA + GAIN (T5-3B) 76.3 81.5 88.5 91.2 73.7 80.0 71.8 77.8
GPT-3.5-turbo (5-shot) 66.6 71.4 82.7 85.3 60.5 66.3 61.9 67.2
Table 4: EM and F1 scores (%) on the hidden test set of GrailQA.
5.2 Implementation Details
Our experiments are done on the machine with an NVIDIA A100 GPU and up to 504GB of RAM. We implement our models utilizing PyTorch Paszke et al. (2019) and Hugging Face [^8] https://huggingface.co/. TIARA+GAIN (T5-3B) takes about 100 hours to train the logical form generator on the synthetic dataset.
Training Question Generator
We fine-tune the T5-base model Raffel et al. (2020) to convert S-expression or triple to natural language questions. We set the beam size to 10, the learning rate to 3e-5, the number of epochs to 10, and the batch size to 8.
Training TIARA
The training of the TIARA model Shu et al. (2022) follows its original settings, including the setting of hyperparameters and the calculation of metrics. Note that Hits@1 on TIARA is obtained by randomly selecting one answer for each question 100 times. Both the schema retriever and generator of TIARA are pre-trained on synthetic data and then fine-tuned on KBQA datasets. Since GraphQuestions has no official training-valid split, we randomly take 200 questions from the original training set as the valid set.
Training the Relation Linking Model
We use the BERT-base-uncased model Devlin et al. (2019) to rank candidate relations for SQB, and the input form is the same as the schema retriever of TIARA. We set the learning rate to 3e-5, the batch size to 256, and the maximum number of epochs to 3 with early stopping.
5.3 Data Augmentation
The statistics of GAIN-synthetic datasets for both logical forms and triples are shown in Table 2 and 3999Details of synthetic data are shown in Appendix  B.. Note that theoretically, the sampling of the GAIN method is not limited to the scale of the synthetic data we use here.
6 Analysis
Model on GraphQuestions F1($\uparrow$) Std($\downarrow$)
GraphQuestions on Freebase 2013-07
UDepLambda Reddy et al. (2017) 17.7 -
PARA4QA Dong et al. (2017) 20.4 -
SPARQA Sun et al. (2020) 21.5 -
BERT + Ranking Gu et al. (2021) 25.0 -
ArcaneQA Gu and Su (2022) 31.8 -
${}^{\clubsuit}$ (T5-base) Shu et al. (2022) 37.9 0.141
KB-BINDER (6) Li et al. (2023) 39.5 -
TIARA + GAIN (T5-base) 45.5 0.153
TIARA + GAIN (T5-3B) 48.7 0.180
GraphQuestions on Freebase 2015-08-09
BERT + Ranking Gu et al. (2021) 27.0 -
ArcaneQA Gu and Su (2022) 34.3 -
${}^{\clubsuit}$ (T5-base) Shu et al. (2022) 41.2 0.157
Pangu (Codex) Gu et al. (2022a) 44.3 -
Pangu (T5-3B) Gu et al. (2022a) 62.2 -
TIARA + GAIN (T5-base) 49.5 0.170
TIARA + GAIN (T5-3B) 53.0 0.200
Table 5: F1 scores (%) and average standard deviation (std) of F1 scores for each set of paraphrases on the test set of GraphQuestions. The setting for Freebase 2015-08-09 is described by Gu and Su (2022). ${}^{\clubsuit}$ denotes our replication results.
Overall Seen Unseen
Model on SimpleQuestions-Balance 1 5 10 1 5 10 1 5 10
HR-BiLSTM Wu et al. (2019) 63.3 - - 93.5 - - 33.0 - -
Adversarial-Adapter Wu et al. (2019) 84.9 - - 92.6 - - 77.1 - -
BERT-base 83.7 95.0 96.9 85.8 95.0 96.0 81.5 95.1 97.8
BERT-base + GAIN 88.4 96.0 97.3 87.8 95.4 96.3 89.1 96.7 98.4
Table 6: Hits@$k$ (1, 5, 10) scores (%) for relation linking on the test set of SimpleQuestions-Balance, including seen and unseen relations.
Std Overall Std I.I.D. Std Compositional Std Zero-shot
Model on GrailQA Valid Set EM($\downarrow$) F1($\downarrow$) EM F1 EM F1 EM F1
TIARA ELF only Shu et al. (2022) 0.066 0.062 0.017 0.016 0.148 0.152 0.198 0.181
TIARA (T5-base) Shu et al. (2022) 0.079 0.066 0.021 0.017 0.211 0.203 0.222 0.181
TIARA + GAIN ELF only 0.063 0.060 0.017 0.017 0.138 0.148 0.187 0.169
TIARA + GAIN (T5-base) 0.077 0.061 0.020 0.016 0.215 0.198 0.218 0.160
TIARA + GAIN (T5-3B) 0.075 0.058 0.020 0.016 0.196 0.180 0.212 0.155
GPT-3.5-turbo (5-shot) 0.093 0.091 0.027 0.023 0.272 0.281 0.251 0.247
Table 7: Average standard deviation of EM and F1 scores for each set of paraphrases on the GrailQA valid set. ELF denotes exemplary logical form Shu et al. (2022).
Schema-Level Generalization
As shown in Tables 4 and 5, the models perform significantly better on i.i.d. than compositional and zero-shot generalization, with the zero-shot generalization being the most challenging. Besides, an increased number of model parameters, combined with richer data from GAIN, significantly enhance the generalization capabilities of T5 models. TIARA+GAIN achieves the highest EM scores, including that on zero-shot scenes. This demonstrates promising ideas for further improving LM generalization capabilities, i.e., the positive effect of synthetic data and parametric scales on training LMs. However, it is important to note that fine-tuned models consistently outperform few-shot learning models, regardless of whether the schema is seen or not. Given the training and inference costs of LLMs, their performance has yet to show any superiority in this semantic parsing task.
Model on WebQSP F1 Hits@1
${}^{\clubsuit}$ (T5-base) Shu et al. (2022) 28.5 27.6
$*{}^{\clubsuit}$ (T5-base) Shu et al. (2022) 33.5 31.5
BERT + Ranking* Gu et al. (2021) 43.0 -
TIARA + GAIN (T5-base) 29.1 28.2
TIARA + GAIN (T5-3B) 29.8 28.7
TIARA* + GAIN (T5-base) 33.9 31.8
TIARA* + GAIN (T5-3B) 34.5 32.3
Table 8: F1 and Hits@1 scores (%) on WebQSP without fine-tuning on it. All models are trained on large-scale GrailQA. * denotes using oracle entity annotations. ${}^{\clubsuit}$ denotes our replication results.
GrailQA GraphQ WebQSP SQB
Train size 44,337 2,381 3,097 75,819
Valid size 6,763 - - 11,141
Test size 13,231 2,395 1,638 21,483
Length 62.96 54.62 35.93 42.16
# of entities 0.903 1.028 1.112 1.000
# of relations 1.358 1.434 1.464 1.000
Similarity between questions and KB items
Entity 0.999 1.000 0.921 0.985
Class 0.547 0.457 - -
Relation 0.470 0.389 0.300 0.779
Unseen ratio (%)
Schema 16.90 86.78 20.44 32.67
Question 54.06 98.25 4.03 49.18
Table 9: KBQA dataset statistics. Length denotes the average number of characters in the question. # of entities/relations denotes the average number of entities/relations contained in the logical form. Unseen Schema is the ratio of schema items in the dev/test set that do not appear in the training set. Unseen Question is the ratio of questions containing unseen schema.
Overall I.I.D. Compositional Zero-shot
Model on GrailQA Valid Set EM F1 EM F1 EM F1 EM F1
BERT + Ranking (Gu et al., 2021) 51.0 58.4 58.6 66.1 40.9 48.1 51.8 59.2
TIARA ELF only Shu et al. (2022) 67.2 72.9 72.8 76.7 55.3 60.7 69.7 76.3
RnG-KBQA Ye et al. (2022) 71.4 76.8 86.7 89.0 61.7 68.9 68.8 74.7
DecAF (FiD-3B) Yu et al. (2022) - 81.4 - 89.7 - 80.1 - 78.4
TIARA (T5-base) Shu et al. (2022) 75.3 81.9 88.4 91.2 66.4 74.8 73.3 80.7
Pangu (T5-3B) Gu et al. (2022a) 75.8 83.4 - - - - - -
TIARA + GAIN ELF only 67.4 73.6 72.7 77.5 54.7 60.5 70.3 77.3
TIARA + GAIN (T5-base) 77.1 83.5 89.0 91.9 68.6 75.5 75.4 83.2
TIARA + GAIN (T5-3B) 77.1 83.8 89.0 92.1 68.8 76.1 75.4 83.4
GPT-3.5-turbo (5-shot) 69.7 74.8 83.0 85.5 58.7 64.6 68.6 74.4
Table 10: EM and F1 scores (%) on the GrailQA valid set. ELF denotes exemplary logical form Shu et al. (2022).
Paraphrase Adaptation
To evaluate the consistency of the model performance across varying expressions of the same semantics, we calculate the standard deviation (std) of EM or F1 for questions within each logical form template and subsequently calculate the average std for all templates in the dev/test set. A lower std signifies greater adaptability of the model to different expressions. For GrailQA, the std of EM and F1 decreases with the application of GAIN or an increase in model size. However, in the case of more challenging GraphQuestions, GAIN significantly improves the F1 but also results in a larger std, as shown in Table 5. It suggests that improving paraphrase adaptation seems more difficult when the baseline model (TIARA, T5-base, with only 37.9% F1) still struggles to address most of the dataset. Consequently, the performance gains observed on the KBQA benchmark may not necessarily reflect a deeper understanding of linguistic complexities, but they could simply render the model more sensitive to specific phrases.
Cross-Dataset Transfer
To emulate an unseen real-world scenario, we evaluate the performance of pre-trained models on the human-curated WebQSP dataset without fine-tuning, as depicted in Table 8. BERT+Ranking Gu et al. (2021) and TIARA+GAIN Shu et al. (2022) are trained on the large-scale GrailQA dataset. We compare these results to the state-of-the-art Pangu Gu et al. (2022a), which is fine-tuned on WebQSP and achieves an F1 score of 79.6%. Although we recognize that GAIN and large models offer few advantages, the performance of these pre-trained models without fine-tuning is considerably lower than Pangu’s. We attribute this to the significant differences between training and test data, as shown in Table 9. The question length, the difficulty of entity/relation linking [^10] Measured by literal similarity: https://anhaidgroup.github.io/py_stringmatching/v0.3.x/PartialRatio., and the proportion of unseen schema vary dramatically across KBQA datasets. These discrepancies arise from the dataset construction process: WebQSP is an annotation of search logs, whereas the remaining datasets are derived from graph search and crowdsourcing. To further enhance robustness in cross-dataset transfer, we believe that better data collection methods are required to obtain diverse and balanced training data. Additionally, the representation of the logical form increases the transfer difficulty, as the S-expression used in GrailQA dataset cannot express all queries in WebQSP.
LLM with In-context Learning
We evaluate the performance of GPT-3.5 with in-context learning on the GrailQA dataset. In the prompt, we provide the task description and the few-shot KB contexts. As illustrated in Table 10, when provided with contexts from the TIARA+GAIN retrievers, GPT-3.5 outperforms the ELF contexts but falls short compared to T5 generators. Among the GPT-3.5 predictions, 79.62% come directly from the substring of the corresponding prompts, achieving an average F1 score of 86.19% for this portion. However, the remaining predictions are not part of their prompts and are entirely new predictions generated by GPT-3.5, with an average F1 score of merely 30.29%. Although a baseline level is attained, these results suggest that GPT-3.5 cannot be accurately grounded to the KB environment when it does not directly utilize the retrievers’ contexts. It also means the severance of natural language pre-training and KB contexts for the LLM, and the faithfulness and controllability of grounding LLMs are not yet guaranteed under the current approach Gu et al. (2022a). To mitigate this problem, alternative paradigms should be explored, such as tool learning Schick et al. (2023) and multi-step planning Liu et al. (2023), which enables more refined access and control over environments.
7 Conclusion
We find that despite the voluminous training corpus available for LMs, incorporating data that adequately captures the intricate complexities of environments such as large-scale KBs is not yet fulfilled. This underscores the importance of focusing on the robustness challenges. Notably, LLMs sometimes blindly follow the provided prompt. It indicates that the existing methodologies for grounding LLMs are yet to prove their efficacy and superiority. Future research issues include collecting more balanced environment-specific corpora and improving the LLM learning paradigm. For the corpus collection problem, our experiments show some potential for data augmentation techniques.
Limitations
For question generation, the verbalization process of the GAIN method relies heavily on large-scale KBQA annotations. The generated questions may be similar to the training data, and overly complex logical forms (e.g., with three and more hops) are difficult to convert into natural language questions. Besides, synthetic data is less diverse and natural than human annotations, though it improves generalization performance.
Ethics Statement
The proposed approach GAIN could be used on any KB for data augmentation. The Freebase Bollacker et al. (2008) used in this work is a KB that has been publicly released and manually reviewed. For uncensored KBs, if harmful information is collected, it could make synthetic data contain harmful information and make LMs generate harmful answers.
Appendix A Example of LLM Prompt
We present an example of an excerpted prompt, which is shown in Figures 2 and 3. In particular, Figure 2 illustrates the task instruction and teaching example segment, while Figure 3 displays the input query segment.
Appendix B Details of Synthetic Data
The graph search process of GAIN is slightly different for logical forms and triples.
Searching Logical Forms
GAIN employs a graph search approach similar to GraphQuestions Su et al. (2016) to collect logical forms derived from graph queries. The graph query construction process consists of four steps: 1) query template construction, 2) aligning template nodes, 3) configuring functions, and 4) verification by execution. Query templates, obtained through random graph searching, contain nodes that represent entity/literal types (not aligned to a value), referred to as template nodes. Each unaligned node in the query template is then aligned with a topic entity or literal to generate multiple aligned graph queries. To synthesize counting, comparative, and superlative questions and enhance the diversity of synthetic data, we add functions like COUNT, ARGMIN/ARGMAX Gu et al. (2021) to graph queries. Since KBQA research typically assumes that questions can be answered by the KB, we execute all resulting graph queries after the above steps and remove those with a null [^11] Null for querying entities and zero for counting. result.
Searching Triples
A single KB triple can be treated as a QA pair, where the head entity and relation together form the query, and the tail entity is the answer. The triple search process consists of two steps: 1) candidate relation selection, and 2) triple sampling. First, arbitrary relations $\mathcal{R}$ are selected from the KB, without being restricted to any particular KBQA dataset. Then, triples are collected from head entities $\mathcal{H}$, where entities in $\mathcal{H}$ are connected to relations in $\mathcal{R}$.
Examples of Synthetic Data
We present some examples of synthetic data in Table 15, where the logical form contains a variety of functions.
Statistics
The statistics for KBQA datasets including our synthetic dataset are shown in Table 16. To calculate the number of domains in Table 2, 3, we take the domain of each Freebase class from its first segment, except for classes starting with “base”, where we take the first two segments, e.g., domain “music” for the class “music.performance_venue”, and domain “base.plants” for the class “base.plants.plant”.
Metrics GrailQA Dev GraphQ Test SQB Dev
BLEU-4 0.347 0.178 0.369
ROUGE-L 0.526 0.411 0.640
Table 13: The performance of the question generator on KBQA datasets. The generator for logical form is evaluated on the GrailQA valid set and GraphQuestions test set. The generator for triple is evaluated on the SQB valid set.
Logical Form (S-expression) Question
1. (AND book.journal (JOIN book.periodical.editorial_staff
(AND (JOIN book.editorial_tenure.editor m.012z2ncg)
(JOIN book.editorial_tenure.title m.02h6676))))
(GrailQA valid set)
2. (AND book.journal (JOIN book.periodical.editorial_staff
(AND (JOIN book.editorial_tenure.editor m.05ws_t6)
(JOIN book.editorial_tenure.title m.02wk2cy))))
(GrailQA valid set)
1. john oliver la gorce was the editor on the editor for what journal?
2. with which journal did don slater serve as editor on the editor in chief?
All four S-expressions are (COUNT (AND book.reviewed_work
(JOIN book.reviewed_work.reviews_of_this_work m.0240y2)))
(GraphQuestions training set)
1. how many works did fresh air review?
2. how many works were reviewed by fresh air in total?
3. what is the total amount of works reviewed by fresh air?
4. fresh air has reviewed how many different works?
Table 14: Examples of paraphrases in GrailQA and GraphQuestions.
Sampled Logical Form Synthetic Question
(COUNT (AND people.profession (JOIN
people.profession.people_with_this_profession m.012d40)))
|entity|m.012d40 jackie chan
how many professions does jackie chan have?
(AND food.beer (le food.beer.original_gravity 1.067^^float)) which beer has an original gravity less than or equal to 1.067?
(AND medicine.manufactured_drug_form (AND (lt
medicine.manufactured_drug_form.size 10.0^^float)
(JOIN medicine.manufactured_drug_form.fda_otc_part m.0h9yt7z)))
|entity|m.0h9yt7z fda otc monograph part 348
which manufactured drug form has part
fda otc monograph part 348 and has a size smaller than 10.0?
(ARGMAX (AND measurement_unit.power_unit
(JOIN measurement_unit.power_unit.measurement_system m.07y37))
measurement_unit.power_unit.power_in_watts)
|entity|m.07y37 us customary units
what is the largest power unit in the us customary units?
(AND music.release (AND (JOIN music.release.engineers m.011mbx12)
(JOIN music.release.label m.0g12fn3)))
|entity|m.011mbx12 raynard glass |m.0g12fn3 hostyle gospel ministries
what musical release is engineered by raynard glass and
labelled hostyle gospel ministries?
Sampled Triple Synthetic Question
D.W. Robertson, Jr. (m.09ggymq),
people.person.place_of_birth,
Washington, D.C. (m.0rh6k)
where was D. W. Robertson, Jr. born
Alfred Chao (m.046cmd8),
computer.operating_system_developer.operating_systems_developed,
pSOS (m.0lscq)
what operating system did Alfred Chao develop?
Table 15: Examples of synthetic data. The logical form is S-expression Gu et al. (2021). The entity label is appended to the logical form.
Datasets #question #class #relation #entity
GrailQA Gu et al. (2021) 64,331 1,534 3,720 32,585
GraphQuestions Su et al. (2016) 5,166 506 596 376
WebQSP Yih et al. (2016) 4,737 408 661 2,593
GAIN-synthetic 127,329 5,078 12,942 46,645
Table 16: Statistics of KBQA datasets and the GAIN-synthetic dataset.
Appendix C Experimental Details
C.1 Question Generation
Training a question generation (QG) model is the second step of GAIN. Because logical forms in the synthetic dataset do not have corresponding human-labeled natural language questions, evaluating QG on the synthetic dataset is difficult. Existing KBQA datasets contain questions written by crowdsourced workers for logical forms or triples, so we evaluate the question generator by generating questions on these datasets and calculate the BLEU-4 Papineni et al. (2002) and ROUGE-L Lin and Och (2004) scores [^12] Calculated by Hugging Face Evaluate. (Table 13).
C.2 Retrieval Performance
The performance of schema retrieval on the GrailQA valid set and GraphQuestions test set is shown in Table 11 and 12, respectively. GAIN significantly improves the performance of schema retrieval on both GrailQA and GraphQuestions.
Function None Count Comparative Superlative
ArcaneQA Gu and Su (2022) 70.8/77.8 62.5/68.2 54.5/75.7 70.5/75.6
RnG-KBQA Ye et al. (2022) 77.5/81.8 73.0/77.5 55.1/76.0 13.8/22.3
TIARA (T5-base) Shu et al. (2022) 77.8/83.1 76.4/81.8 57.4/81.4 58.7/69.0
TIARA + GAIN ELF only 76.8/81.7 73.9/80.0 0.0/25.3 0.0/8.3
TIARA + GAIN (T5-base) 78.6/84.6 77.7/83.0 61.7/82.3 69.9/73.2
TIARA + GAIN (T5-3B) 78.5/84.8 77.3/82.5 63.0/84.5 70.7/74.1
GPT-3.5-turbo (5-shot) 74.1/78.0 66.8/70.5 38.3/60.5 43.9/52.3
# of relations 1 2 3 4
RnG-KBQA Ye et al. (2022) 75.7/79.3 65.3/74.7 28.6/44.5 100.0/100.0
TIARA (T5-base) Shu et al. (2022) 81.2/85.6 64.7/75.8 29.3/48.5 50.0/83.3
TIARA + GAIN ELF only 74.0/77.8 56.6/67.9 9.9/31.0 0.0/33.3
TIARA + GAIN (T5-base) 82.4/87.2 67.0/78.0 38.9/49.8 50.0/83.3
TIARA + GAIN (T5-3B) 82.0/87.2 68.8/79.0 37.5/51.3 50.0/83.3
GPT-3.5-turbo (5-shot) 75.0/78.6 61.9/69.6 19.8/36.6 50.0/50.0
# of entities 0 1 2
RnG-KBQA Ye et al. (2022) 58.5/63.6 75.4/79.9 55.6/73.5
TIARA (T5-base) Shu et al. (2022) 77.5/83.1 76.6/82.6 49.9/68.0
TIARA + GAIN ELF only 42.8/47.0 74.2/79.9 47.4/67.6
TIARA + GAIN (T5-base) 82.2/86.5 77.6/83.7 53.4/71.4
TIARA + GAIN (T5-3B) 82.0/86.6 77.6/84.0 55.6/73.0
GPT-3.5-turbo (5-shot) 66.8/72.8 71.9/76.1 48.8/62.1
Table 17: EM and F1 scores (%) for different types of logical forms on the GrailQA valid set. None denotes no function. # of relations/entities denotes the number of relations/entities in the S-expression. ELF denotes exemplary logical form Shu et al. (2022).
C.3 Performance on Various Logical Forms
To show how TIARA+GAIN performs on different types of logical forms, we compare it with previous KBQA models on the GrailQA valid set, as shown in Table 17. TIARA+GAIN improves performance in nearly all these scenarios compared to TIARA.
C.4 Error Analysis
To analyze the QA errors of TIARA+GAIN (T5-3B), we randomly sample 50 questions where predicted logical forms are not the same as the ground truth in the GrailQA valid set. We follow Shu et al. (2022) in their classification of errors. Entity linking errors (missing or redundant entities), syntactic errors (generating the wrong logical form structure), semantic errors (generating the wrong schema item when correct contexts are provided), false negatives (flawed annotations), and miscellaneous (e.g., ambiguity in the question) account for 48%, 26%, 16%, 4%, and 6%, respectively. For entity linking errors, 62.5% of them are from the zero-shot level. For syntactic errors, the number of errors from zero-shot and compositional levels is almost the same. It means that entity linking for zero-shot domains and logical form generation for complex structures remain significant challenges.
Given a question and Freebase contexts, write a logical form that answers the question. Question: dark sun: wake of the ravager was designed by what video game designer? Candidate entities: (A) [Dark Sun: Wake of the Ravager] Exemplary Logical Forms: (A) (AND cvg.cvg_designer (JOIN cvg.cvg_designer.games_designed [Dark Sun: Wake of the Ravager])) (B) (AND cvg.cvg_designer (JOIN (R cvg.computer_videogame.designers) [Dark Sun: Wake of the Ravager])) (C) (AND cvg.computer_videogame (JOIN (R cvg.cvg_designer.games_designed) (JOIN cvg.cvg_designer.games_designed [Dark Sun: Wake of the Ravager]))) (D) (AND cvg.computer_videogame (JOIN cvg.computer_videogame.designers (JOIN cvg.cvg_designer.games_designed [Dark Sun: Wake of the Ravager]))) (E) (AND base.wikipedia_infobox.video_game (JOIN base.wikipedia_infobox.video_game.developer (JOIN cvg.cvg_designer.games_designed [Dark Sun: Wake of the Ravager]))) Candidate classes: (A) cvg.cvg_designer (B) cvg.game_performance (C) cvg.musical_game (D) cvg.game_character (E) cvg.computer_game_engine_developer (F) cvg.computer_videogame (G) cvg.computer_game_performance_type (H) cvg.game_version (I) cvg.computer_game_subject (J) cvg.computer_game_evaluation Candidate relations: (A) cvg.cvg_designer.games_designed (B) cvg.computer_videogame.designers (C) cvg.computer_videogame.prequel (D) cvg.computer_videogame.sequel (E) cvg.computer_videogame.mods (F) cvg.computer_videogame.expansions (G) cvg.computer_videogame.developer (H) cvg.computer_videogame.characters (I) cvg.game_version.game (J) cvg.computer_game_mod.game_modded Prediction: (AND cvg.cvg_designer (JOIN cvg.cvg_designer.games_designed [Dark Sun: Wake of the Ravager]))
Figure 2: Example of LLM prompt (part 1): The task instruction and $k$ teaching examples (only one is shown because of the length) with questions and their contexts.
Question: worldofwarcraft is the creation of which video game designer? Candidate entities: (A) [worldofwarcraft] Exemplary Logical Forms: (A) (AND cvg.cvg_designer (JOIN cvg.cvg_designer.games_designed [worldofwarcraft])) (B) (AND cvg.cvg_designer (JOIN (R cvg.computer_videogame.designers) [worldofwarcraft])) (C) (AND cvg.cvg_designer (JOIN cvg.cvg_designer.games_designed (JOIN cvg.computer_game_expansion.expansion_for [worldofwarcraft]))) (D) (AND cvg.cvg_designer (JOIN (R cvg.computer_videogame.designers) (JOIN cvg.computer_game_expansion.expansion_for [worldofwarcraft]))) (E) (AND cvg.cvg_designer (JOIN (R cvg.computer_videogame.designers) (JOIN (R cvg.computer_videogame.expansions) [worldofwarcraft]))) Candidate classes: (A) games.game_designer (B) cvg.cvg_designer (C) amusement_parks.ride_designer (D) cvg.cvg_developer (E) cvg.computer_videogame (F) cvg.computer_game_engine_developer (G) cvg.computer_game_engine (H) cvg.computer_game_mod (I) cvg.game_performance (J) cvg.musical_game Candidate relations: (A) cvg.computer_videogame.designers (B) cvg.cvg_designer.games_designed (C) games.game_designer.games_designed (D) games.game.designer (E) cvg.computer_videogame.developer (F) cvg.cvg_developer.games_developed (G) cvg.computer_game_engine.developer (H) cvg.computer_videogame.expansions (I) cvg.computer_videogame.publisher (J) cvg.game_version.developer Prediction:
Figure 3: Example of LLM prompt (part 2): After $k$ teaching examples, the input question and its contexts is given. The golden prediction is (AND cvg.cvg_designer (JOIN cvg.cvg_designer.games_designed [worldofwarcraft])).