### 09 Feb 2024 v3

# Distribution Shifts Are Bottlenecks: Extensive Evaluation for Grounding Language Models to Knowledge Bases

Abstract
Grounding language models (LMs) to knowledge bases (KBs) helps to obtain rich and accurate facts. However, it remains challenging because of the enormous size, complex structure, and partial observability of KBs. One reason is that current benchmarks fail to reflect robustness challenges and fairly evaluate models. This paper analyzes whether these robustness challenges arise from distribution shifts, including environmental, linguistic, and modal aspects. This affects the ability of LMs to cope with unseen schema, adapt to language variations, and perform few-shot learning. Thus, the paper proposes extensive evaluation protocols and conducts experiments to demonstrate that, despite utilizing our proposed data augmentation method, both advanced small and large language models exhibit poor robustness in these aspects. We conclude that current LMs are too fragile to navigate in complex environments due to distribution shifts. This underscores the need for future research focusing on data collection, evaluation protocols, and learning paradigms. [^1] Code and data are available at https://github.com/yhshu/Distribution-Shifts-for-KBQA.
Distribution Shifts Are Bottlenecks: Extensive Evaluation for
Grounding Language Models to Knowledge Bases
Yiheng Shuthanks: * Work performed when the author was graduate student at Nanjing University. The Ohio State University shu.251@osu.edu                        Zhiwei Yu Microsoft zhiwyu@microsoft.com
1 Introduction
Language models (LMs), such as BERT Devlin et al. (2019), T5 Raffel et al. (2020), and the GPT series Ouyang et al. (2022); OpenAI (2023), have demonstrated impressive capabilities in understanding and generating languages, highlighting the potential for artificial general intelligence (AGI). However, a major obstacle to achieving this goal is that LMs mainly built on natural languages are not yet well-grounded to real-world environments, such as knowledge base (KB), an environment of enormous size, complex structure, and only partially observable to LM.
Figure 1: Distribution shifts during grounding in the case of KBQA. Training and inference using LM face completely different data distributions. We call for these shifts to be reflected in the evaluation protocols.
Though LMs are highly skilled at natural language question answering (QA) today, the task of Knowledge Base Question Answering (KBQA) aims to parse natural language queries into formal queries on KBs, such as Freebase Bollacker et al. (2008) and Wikidata Vrandecic and Krötzsch (2014). The significance of this task lies in building language agents on complex environments Su (2023), rather than merely recalling answers from the LM’s stored knowledge.
Now, numerous LM-driven models Das et al. (2021); Hu et al. (2022) continue to achieve higher F1/Hits@1 scores on KBQA benchmarks. However, achieving higher scores does not necessarily guarantee the development of robust and dependable models. We still need to ask whether such improvements apply to extensive scenarios, as benchmarks almost always create questions via crowdsourcing and evaluate with simplistic metrics (Table 1). These benchmarks may not fully represent the diverse scenarios encountered in real-world applications, which raises concerns about the robustness of LM-driven models. Thus, our research aims to bridge this gap by exploring the limitations of current KBQA benchmarks and proposing more comprehensive evaluation protocols.
To achieve this goal, we need to grasp the key factors in robustness. For modern deep learning systems, the amount of training data could be extremely rich, but robustness is closely related to data distribution Hendrycks et al. (2020). In the general area of natural language processing (NLP), large-scale corpora have been collected and used for effective training Touvron et al. (2023). However, real-world environments are rarely so accommodating, e.g., large KBs contain complex structures and schema items, and building a large-scale and representative corpus is quite challenging. The problem of inconsistent data distribution during training and inference, i.e., distribution shifts as shown in Figure 1, may negatively impact the performance and robustness of grounded LMs.
This paper analyzes the distribution shifts from multiple aspects to understand both KBQA benchmarking and LM grounding better. We review existing works and identify several challenges. 1) Environmental aspect: generalization to unseen domains at the schema [^2] Schema denotes rdfs:Class (class) and rdf:Property (relation) here. level Gu et al. (2021); 2) Linguistic aspect: adaptation to paraphrases featuring diverse language variations Su et al. (2016); 3) Integrated aspect: transferability across datasets with both novel schema items and linguistic styles Cao et al. (2022b); 4) Modal aspect: few-shot learning ability of large language models (LLMs) for the unfamiliar KB modality Li et al. (2023). These challenges guide us to build more rigorous evaluation protocols for multiple aspects (§4).
To measure the impact of distribution shifts, we conduct extensive experiments under the proposed evaluation protocols. To present a fair evaluation under such rigorous protocols, we implement a data augmentation method for evaluated LMs and a retrieval augmentation method for evaluated LLMs (§5.1). Our findings reveal that even when employed with such methods and the highest EM scores are achieved on the GrailQA benchmark Gu et al. (2021), advanced small and large LMs still fall short of effectively tackling the majority of these challenges. A striking example is the large difference between the best practice without WebQSP Yih et al. (2016) fine-tuning (F1 43.0%) compared to the fine-tuned state-of-the-art (F1 79.6%) (§6), suggesting the weak robustness of LM-driven KBQA models on an unseen dataset. Such negative observations highlight an urgent need for future research in data collection methodologies and LM learning paradigms. Meanwhile, we expect our evaluation protocols to provide a reference for future benchmark construction, developing metrics that consider robustness.
Our contributions include 1) A systematic analysis of the robustness challenges of grounded LMs and a strong advocacy of a multi-aspect evaluation protocol for KBQA. 2) Extensive experiments quantitatively unveil the existence of these challenges and the vulnerability of both small and large LMs. 3) Insights into improving the robustness of grounded LMs, including data collection and learning paradigms.
2 Related Work
Compared to existing research, the necessity of this study is threefold. First, while LMs are commonly evaluated on natural language tasks Hupkes et al. (2022); Patel et al. (2022), the complexity increases significantly when these models are applied to environments like KBs, where data is structured rather than purely unstructured natural language Liu et al. (2023). Second, the evaluation protocols used in KBQA benchmarks tend to be uniform, leading to an insufficient evaluation of model robustness Gu et al. (2021). Finally, recent reviews in this field Lan et al. (2022); Gu et al. (2022b) have largely overlooked the advancements in the development and application of LMs, particularly LLMs.
Benchmark KB Size LF Generalization Paraphrases Metrics
WebQuestions Berant et al. (2013) Freebase 5,810 N/A i.i.d. F1
SimpleQuestions Bordes et al. (2015) Freebase 108,442 N/A i.i.d. Acc
WebQuestionsSP Yih et al. (2016) Freebase 4,737 SPARQL i.i.d. F1
GraphQuestions Su et al. (2016) Freebase 5,166 Graph query comp.+zero F1
LC-QuAD Trivedi et al. (2017) DBpedia 5,000 SPARQL i.i.d. F1
CWQ Talmor and Berant (2018) Freebase 34,689 SPARQL i.i.d. Hits
LC-QuAD 2.0 Dubey et al. (2019) Wikidata 30,000 SPARQL i.i.d. F1
SQB Wu et al. (2019) Freebase 108,443 N/A i.i.d.+zero Acc
CFQ Keysers et al. (2020) Freebase 239,357 SPARQL comp Acc
GrailQA Gu et al. (2021) Freebase 64,331 S-expression i.i.d.+comp.+zero EM, F1
KQA Pro Cao et al. (2022a) Wikidata 117,970 KoPL i.i.d. Acc
QALD series Perevalov et al. (2022) DBpedia 558 SPARQL comp. F1
Table 1: Selected KBQA benchmarks. LF denotes logical forms. Generalization settings follow Gu et al. (2021). i.i.d. denotes that the schema distribution in the test set is the same as the training set. comp. and zero denote compositional and zero-shot generalization, respectively. Paraphrases are questions containing the same semantics (machine-generated paraphrases are not included). Acc denotes accuracy. EM denotes exact match.
3 Challenges from Distribution Shifts
In this paper, the robustness of a KBQA model refers to its ability to adapt to various natural language inputs and maintain consistent performance when data distribution shifts. Due to the distribution shifts between the training corpus of LMs and KB environments, grounding LMs face robustness challenges from environmental, linguistic, and modal aspects.
3.1 Environmental Aspect
A primary environmental challenge is schema-level generalization. The RDF Schema provides a data-modeling vocabulary crucial for querying a KB. Table 1 indicates most KBQA benchmarks assume a consistent schema distribution between training and testing. However, this often does not hold in large KBs with numerous schema items. Few benchmarks, like the reconfigured SimpleQuestions-Balance dataset Wu et al. (2019) and GrailQA Gu et al. (2021), address non-i.i.d. schema items and varying levels of schema-level generalization. GraphQuestions Su et al. (2016) provides a stringent test with seldom-seen schema items in training. Given real-world non-i.i.d. complexities, these datasets better represent the practical generalization challenges. Yet, despite progress Shu et al. (2022); Gu et al. (2022a), compositional and zero-shot generalization are far from solved. We stress that this challenge applies to the overall KBQA semantic parsing process, especially to important subtasks such as relation linking.
3.2 Linguistic Aspect
Various natural language expressions make question understanding challenging for KBQA models. One common way this variety shows up is through paraphrasing. In this paper, a paraphrase set denotes different ways to express the same logical form, as illustrated in Table 14. Paraphrase adaptation is an intuitive form of the ability to comprehend variable language expressions. It could be measured by whether a model is able to accurately answer paraphrased questions that this model has already answered correctly before. Unfortunately, as shown in Table 1, many KBQA benchmarks do not account for paraphrasing with only one utterance for each logical form. Exceptionally, some datasets Su et al. (2016); Dubey et al. (2019); Gu et al. (2021) are based on automatically generated logical forms and include multiple natural language expressions for the same logical form (template). These data characteristics highlight the difficulties in adapting to paraphrased questions.
3.3 Integrated Aspect
Evaluating KBQA benchmarks often hinges on a single dataset, thereby complicating the task of ascertaining the model performance consistency across novel scenarios. This form of robustness, termed as cross-dataset transfer in this paper, combines both the environmental and linguistic aspects discussed earlier and is more difficult to achieve. This is because construction methods vary across datasets, as do schema distributions and natural language expressions. Specifically, KBQA dataset construction generally falls into two distinct categories: 1) Graph Search and Crowdsourcing: in this approach, logical forms or triples are initially extracted from a KB, where structures or operators of logical form are usually finite. Subsequently, they are converted into natural language utterances through crowdsourcing methods Bordes et al. (2015); Trivedi et al. (2017). 2) Human Curation and Parsing: logical forms are labeled directly from human-provided utterances Berant et al. (2013); Perevalov et al. (2022). Existing works Gu et al. (2021); Cao et al. (2022b) suggest that models pre-trained on large-scale datasets can adapt reasonably well to other target datasets, such as WebQSP Yih et al. (2016). However, the necessity for fine-tuning these pre-trained models on the intended target dataset remains imperative for achieving optimal performance. Despite the advantages offered by pre-training on expansive KBQA datasets, models still encounter challenges in transferring directly to previously unseen target datasets while sustaining high performance.
3.4 Modal Aspect
Aside from considering environmental and linguistic factors, focusing on the modal aspect is also crucial. Recently, LLMs like GPT series OpenAI (2023) have demonstrated exceptional capabilities across a variety of tasks, outperforming smaller yet potent LMs such as BERT Devlin et al. (2019) and T5 Raffel et al. (2020). Despite these advancements, these LLMs face substantial challenges when interacting with environments. One notable issue is they predominantly rely on an in-context learning paradigm as opposed to fine-tuning, as a trade-off between computational cost and model efficiency. In comparison to fine-tuning, in-context learning offers the advantage of reduced training costs but at the expense of being forced to reason over the unfamiliar modality. Distribution shifts between natural language pre-training and reasoning over structured knowledge contexts could lead to poor performance. For instance, a discernible performance gap exists between KBQA models that employ in-context learning with Codex Chen et al. (2021a) and those built on fine-tuned LMs Gu et al. (2022a); Li et al. (2023). However, the empirical specifics of this difference are not yet clear, leaving us with an inadequate understanding of the limitations of in-context learning and ways to improve grounding with LLMs.
4 Evaluation Protocols
Regarding these challenges, we introduce extensive protocols for evaluating LMs in several aspects overlooked by current benchmarks.
4.1 Evaluating Environmental Aspect
To set the environmental schema-level generalization scenario, we use GrailQA Gu et al. (2021) and GraphQuestions Su et al. (2016) datasets. GrailQA contains three generalization levels: i.i.d. (25%), compositional (25%), and zero-shot (50%). GraphQuestions has no seen relations in the test set. We also use SimpleQuestions-Balance (SQB) Wu et al. (2019) for the relation linking task (an important KBQA sub-task), where 50% of the samples contain unseen relations.
4.2 Evaluating Linguistic Aspect
To set a paraphrase adaptation scenario, we use GrailQA Gu et al. (2021) and GraphQuestions Su et al. (2016) datasets. To evaluate adaptability to paraphrases (§3.2), we propose a new metric, the standard deviation (std) of EM/F1 for questions of each logical form template. As shown in Equation 1, suppose there are $n$ sets of paraphrases in the dataset, each set of paraphrases corresponds to a logical form template with $m$ natural language expressions, and the F1 score obtained by the KBQA model on the $j$-th question of the $i$-th set of paraphrases is $F1_{i,j}$. The metric $Std_{F1}$ first calculates the standard deviation of the F1 scores obtained by the model on the $m$ questions for each set of paraphrases and then calculates the average of the $n$ standard deviations. This metric is used to measure the robustness of the model to different representations of the same semantics, i.e., whether it can cope with diverse natural language expressions. A lower standard deviation indicates that the model is more adaptive to different expressions. $Std_{EM}$ is calculated in the same way.
$Std_{F1}=\frac{1}{n}\sum_{i=1}^{n}\sqrt{\left(\frac{\sum_{j=1}^{m}(F1_{i,j}-\bar{F1_{i}})^{2}}{m}\right)}$ (1)
4.3 Evaluating Integrated Aspect
To emulate a real-world scenario with unknown schema and linguistic distribution for the integrated aspect, we evaluate the performance of pre-trained models on the unseen human-curated WebQSP Yih et al. (2016) dataset, where the questions are derived from search logs and are more realistic. This is a scenario where the distribution changes significantly, as most benchmarks create questions by sampling logical forms and annotating natural language questions via crowdsourcing Lan et al. (2022), where questions are confined to sampled logical forms.
4.4 Evaluating Modal Aspect
To test the capability of the in-context learning paradigm to inference from the KB modality rather than from pure texts, we retrieve structured KB contexts as prompt to evaluate the LLM without particular fine-tuning on KB (§5.1.2). KBs are structured and expansive, but in this case, LLM can only encode a portion of linearized KB contexts.
5 Experiments
5.1 Augmentation Approach
To ensure fair evaluation and fully harness the capabilities of LMs under our extensive and rigorous evaluation protocols, we suggest two strategies to counteract distribution shifts: data augmentation and retrieval augmentation.
5.1.1 Data Augmentation for LMs
Off-the-shelf datasets of limited size may make LM easily overfitted and not adaptable to large KBs. To address the problem that many domains in the KB are often not collected as training data, we propose a data augmentation method named Graph seArch and questIon generatioN (GAIN). Some data augmentation or question generation models Bi et al. (2020); Guo et al. (2022) are only evaluated by the qualify of generated sentences rather than evaluated by the QA task, but GAIN directly serves our KBQA evaluations. Besides, compared to the previous work Hu et al. (2019) that only considers generating questions for triples to help KBQA, GAIN applies to KBQA corresponding to both logical forms and triples. GAIN scales data volume and distribution through four steps: 1) Graph search: Sampling logical forms or triples from arbitrary domains in the KB without being restricted to any particular KBQA dataset. 2) Training question generator: learning to convert logical forms or triples into natural language questions on existing KBQA datasets. 3) Verbalization: Using the question generator from step 2 to verbalize sampled logical forms or triples from step 1, thus creating synthetic questions. 4) Training data expansion: Before fine-tuning any neural models on KBQA datasets, GAIN-synthetic data can be used to train these models or to expand the corpus of in-context samples for LLMs. That is, as a data augmentation method, GAIN is not a KBQA model, but it is used to augment a base KBQA model.
5.1.2 Retrieval Augmentation for LLMs
As the trade-off between cost and effectiveness, we experiment with the prevalent in-context learning paradigm but attempt to improve the quality of in-context samples. We use advanced retrieval methods based on smaller LMs as plug-ins to augment the LLM, similar to the SuperICL approach Xu et al. (2023). Specifically, our steps to generate an LLM prompt for each question include the following. 1) Given an input question, we retrieve $k$ questions ($k$-shot) with BM25 Robertson et al. (2009) from the corpus (the combination of KBQA training set and the GAIN-synthetic dataset). 2) The role of retrieval augmentation for KB environments has been shown by fine-tuned LMs Shu et al. (2022). To assist with grounding LLM, we retrieve KB contexts with off-the-shelf retrievers for $k$ samples and the input question. [^3] The prompt example is demonstrated in Appendix A.
5.2 Setup
Data
All experiments use S-expression Gu et al. (2021) as the logical form due to its clear and concise structure. Entity linking results are taken from TIARA Shu et al. (2022) for GrailQA and WebQSP, and ArcaneQA Gu and Su (2022) for GraphQuestions, because of their public availability and performance.
Model
Compared models are mainly selected from the leaderboard. [^4] https://dki-lab.github.io/GrailQA/ The performances are taken from their papers. For the relation linking task on SQB, we use BERT Devlin et al. (2019) as the base model for GAIN. For KBQA tasks, we use the open-source advanced model TIARA Shu et al. (2022) as the base model for GAIN, due to its strong performance on zero-shot schema items. [^5] Pangu Gu et al. (2022a) also uses entity linking results from TIARA. TIARA is composed of multi-grained retrievers and a generator, with the retrievers providing KB contexts [^6] Entities, exemplary logical forms, and schema items are retrieved. for the generator. The term “TIARA+GAIN” represents a model (both the retrievers and the generator) that is first tuned using GAIN synthetic data and subsequently fine-tuned on a target dataset. For LLM evaluation in the modal aspect, we use the gpt-3.5-turbo-0613777https://platform.openai.com/docs/models model, and the few-shot contexts are retrieved from the combination of GrailQA training set and synthetic dataset using the TIARA+GAIN retrievers.
Metrics
Following previous works, we use Exact Match (EM), F1, and Hits@1 to measure the performance of KBQA models. We also use the std of EM/F1 to measure the adaptability to paraphrases (§4.2).
Overall I.I.D. Compositional Zero-shot
Model on GrailQA Test Set EM F1 EM F1 EM F1 EM F1
Fine-tuned Models
BERT + Ranking (Gu et al., 2021) 50.6 58.0 59.9 67.0 45.5 53.9 48.6 55.7
RnG-KBQA (Ye et al., 2022) 68.8 74.4 86.2 89.0 63.8 71.2 63.0 69.2
TIARA (T5-base) (Shu et al., 2022) 73.0 78.5 87.8 90.6 69.2 76.5 68.0 73.9
DecAF (FiD-3B) (Yu et al., 2022) 68.4 78.8 84.8 89.9 73.4 81.8 58.6 72.3
Pangu (BERT-base) (Gu et al., 2022a) 73.7 79.9 82.6 87.1 74.9 81.2 69.1 76.1
Pangu (T5-large) (Gu et al., 2022a) 74.8 81.4 82.5 87.3 75.2 82.2 71.0 78.4
Pangu (T5-3B) (Gu et al., 2022a) 75.4 81.7 84.4 88.8 74.6 81.5 71.6 78.5
Codex-driven Models
KB-BINDER (6)-R  (Li et al., 2023) 53.2 58.5 72.5 77.4 51.8 58.3 45.0 49.9
Pangu (Codex) (Gu et al., 2022a) 56.4 65.0 67.5 73.7 58.2 64.9 50.7 61.1
GAIN-augmented Models
TIARA + GAIN (T5-base) 75.1 80.6 88.3 91.0 73.0 79.6 69.9 76.4
TIARA + GAIN (T5-3B) 76.3 81.5 88.5 91.2 73.7 80.0 71.8 77.8
GPT-3.5-turbo (5-shot) 66.6 71.4 82.7 85.3 60.5 66.3 61.9 67.2
Table 2: EM and F1 scores (%) on the hidden test set of GrailQA.
5.3 Implementation Details
We use a machine with an NVIDIA A100 GPU and up to 504GB of RAM. Models are implemented by PyTorch Paszke et al. (2019) and Hugging Face. [^8] https://huggingface.co/ TIARA+GAIN (T5-3B) takes about 100 hours to train the logical form generator on the synthetic dataset.
Model Training
1) For question generation, we fine-tune the T5-base model Raffel et al. (2020) to convert S-expression or triple to natural language questions. We set the beam size to 10, the learning rate to 3e-5, the number of epochs to 10, and the batch size to 8. 2) The training of the TIARA model Shu et al. (2022) follows its original settings, including the setting of hyperparameters and the calculation of metrics. Note that Hits@1 on TIARA is obtained by randomly selecting one answer for each question 100 times. Both the schema retriever and generator of TIARA are pre-trained on synthetic data and then fine-tuned on KBQA datasets. Since GraphQuestions has no official training-valid split, we randomly take 200 questions from the original training set as the valid set. 3) We use BERT-base-uncased Devlin et al. (2019) to rank candidate relations for SQB, and the input form is the same as the schema retriever of TIARA. We set the learning rate to 3e-5, the batch size to 256, and the max number of epochs to 3 with early stopping.
Data Augmentation
The statistics of GAIN-synthetic datasets for both logical forms and triples are shown in Table 11 and 12.999Details of synthetic data are shown in Appendix B. Note that the sampling of the GAIN method is not limited to the scale of the synthetic data we use here.
6 Analysis
Model on GraphQuestions F1($\uparrow$) Std($\downarrow$)
GraphQuestions on Freebase 2013-07
UDepLambda Reddy et al. (2017) 17.7 -
PARA4QA Dong et al. (2017) 20.4 -
SPARQA Sun et al. (2020) 21.5 -
BERT + Ranking Gu et al. (2021) 25.0 -
ArcaneQA Gu and Su (2022) 31.8 -
${}^{\clubsuit}$ (T5-base) Shu et al. (2022) 37.9 0.141
KB-BINDER (6) Li et al. (2023) 39.5 -
TIARA + GAIN (T5-base) 45.5 0.153
TIARA + GAIN (T5-3B) 48.7 0.180
GraphQuestions on Freebase 2015-08-09
BERT + Ranking Gu et al. (2021) 27.0 -
ArcaneQA Gu and Su (2022) 34.3 -
${}^{\clubsuit}$ (T5-base) Shu et al. (2022) 41.2 0.157
Pangu (Codex) Gu et al. (2022a) 44.3 -
Pangu (T5-3B) Gu et al. (2022a) 62.2 -
TIARA + GAIN (T5-base) 49.5 0.170
TIARA + GAIN (T5-3B) 53.0 0.200
Table 3: F1 scores (%) and average standard deviation (std) of F1 scores for each set of paraphrases on the test set of GraphQuestions. The setting for Freebase 2015-08-09 is described by Gu and Su (2022). ${}^{\clubsuit}$ denotes our replication results.
Overall Seen Unseen
Model on SimpleQuestions-Balance 1 5 10 1 5 10 1 5 10
HR-BiLSTM Wu et al. (2019) 63.3 - - 93.5 - - 33.0 - -
Adversarial-Adapter Wu et al. (2019) 84.9 - - 92.6 - - 77.1 - -
BERT-base 83.7 95.0 96.9 85.8 95.0 96.0 81.5 95.1 97.8
BERT-base + GAIN 88.4 96.0 97.3 87.8 95.4 96.3 89.1 96.7 98.4
Table 4: Hits@$k$ (1, 5, 10) scores (%) for relation linking on the test set of SimpleQuestions-Balance, including seen and unseen relations.
Std Overall Std I.I.D. Std Compositional Std Zero-shot
Model on GrailQA Valid Set EM($\downarrow$) F1($\downarrow$) EM F1 EM F1 EM F1
TIARA (T5-base) Shu et al. (2022) 0.079 0.066 0.021 0.017 0.211 0.203 0.222 0.181
TIARA + GAIN (T5-base) 0.077 0.061 0.020 0.016 0.215 0.198 0.218 0.160
TIARA + GAIN (T5-3B) 0.075 0.058 0.020 0.016 0.196 0.180 0.212 0.155
GPT-3.5-turbo (5-shot) 0.093 0.091 0.027 0.023 0.272 0.281 0.251 0.247
Table 5: Average standard deviation of EM and F1 scores for each set of paraphrases on the GrailQA valid set.
We report and analyze the experimental results in this section for each aspect.
6.1 Analysis of Environmental Aspect
Effectiveness of Synthesis and Scaling Up
As shown in Tables 2 and 3, the models perform significantly better on i.i.d. than compositional and zero-shot generalization, with the zero-shot partition being the most challenging. TIARA+GAIN (T5-base) improves 2.5 zero-shot F1 points compared to TIARA (T5-base). Besides, an increased number of model parameters, combined with richer data from GAIN, significantly enhance the generalization capabilities of T5 models. TIARA+GAIN (T5-3B) further improves 1.4 zero-shot F1 points compared to its T5-base version. TIARA+GAIN achieves the highest EM scores, including that on zero-shot scenes. It demonstrates promising ideas for further improving LM generalization capabilities, i.e., the positive effect of synthetic data and parametric scales on training LMs.
Fine-tuning Better Than Few-shot Learning
However, it is important to note that fine-tuned models consistently outperform few-shot learning models, regardless of whether the schema is seen or not. Given the training and inference costs of LLMs, their performance has yet to show any superiority in this task.
6.2 Analysis of Linguistic Aspect
Improvements Are Linguistic Biased
We calculate the standard deviation (std) of EM or F1 in the dev/test set, as shown in Equation 1. For GrailQA, the std of EM and F1 decreases with the application of GAIN or an increase in model size, i.e., F1/EM and the std of F1/EM are both better, as shown in Table 2 and 5. However, in the case of more challenging GraphQuestions, GAIN significantly improves the F1 by 8.3 points but also results in a larger std (0.170 compared to 0.157), as shown in Table 3. It suggests that improving paraphrase adaptation using GAIN is more difficult when the base model (TIARA, T5-base, with only 37.9% F1) still struggles to address most of the dataset. Consequently, the performance gains observed on the KBQA benchmark may not necessarily reflect a deeper understanding of linguistic complexities, but they could simply render the model more sensitive to specific phrases. Strategies for deeper decomposition and understanding Hu et al. (2021); Huang et al. (2023) of the questions may be needed to mitigate this challenge.
Model on WebQSP F1 Hits@1
${}^{\clubsuit}$ (T5-base) Shu et al. (2022) 28.5 27.6
$*{}^{\clubsuit}$ (T5-base) Shu et al. (2022) 33.5 31.5
BERT + Ranking* Gu et al. (2021) 43.0 -
TIARA + GAIN (T5-base) 29.1 28.2
TIARA + GAIN (T5-3B) 29.8 28.7
TIARA* + GAIN (T5-base) 33.9 31.8
TIARA* + GAIN (T5-3B) 34.5 32.3
Table 6: F1 and Hits@1 scores (%) on WebQSP without fine-tuning on it. All models are trained on large-scale GrailQA. * denotes using oracle entity annotations. ${}^{\clubsuit}$ denotes our replication results.
Overall I.I.D. Compositional Zero-shot
Model on GrailQA Valid Set EM F1 EM F1 EM F1 EM F1
BERT + Ranking (Gu et al., 2021) 51.0 58.4 58.6 66.1 40.9 48.1 51.8 59.2
TIARA ELF only Shu et al. (2022) 67.2 72.9 72.8 76.7 55.3 60.7 69.7 76.3
RnG-KBQA Ye et al. (2022) 71.4 76.8 86.7 89.0 61.7 68.9 68.8 74.7
DecAF (FiD-3B) Yu et al. (2022) - 81.4 - 89.7 - 80.1 - 78.4
TIARA (T5-base) Shu et al. (2022) 75.3 81.9 88.4 91.2 66.4 74.8 73.3 80.7
Pangu (T5-3B) Gu et al. (2022a) 75.8 83.4 - - - - - -
TIARA + GAIN (T5-base) 77.1 83.5 89.0 91.9 68.6 75.5 75.4 83.2
TIARA + GAIN (T5-3B) 77.1 83.8 89.0 92.1 68.8 76.1 75.4 83.4
GPT-3.5-turbo (5-shot) 69.7 74.8 83.0 85.5 58.7 64.6 68.6 74.4
Table 7: EM and F1 scores (%) on the GrailQA valid set. ELF denotes exemplary logical form Shu et al. (2022).
6.3 Analysis of Integrated Aspect
Hard Transfer Across Datasets
We evaluate the performance of pre-trained models on the human-curated WebQSP dataset without fine-tuning, as shown in Table 6. BERT+Ranking Gu et al. (2021) and TIARA+GAIN Shu et al. (2022) are trained on the large-scale GrailQA dataset. We compare these results to the state-of-the-art Pangu Gu et al. (2022a), which is fine-tuned on WebQSP and achieves an F1 score of 79.6% Although we recognize that GAIN and large models offer few advantages, the performance of these pre-trained models without fine-tuning is considerably lower than Pangu’s.
Causes from Data Collection
We attribute this to the significant differences between training and test data, as shown in Table 8. The question length, the difficulty of entity/relation linking [^10] Measured by literal similarity: https://anhaidgroup.github.io/py_stringmatching/v0.3.x/PartialRatio., and the proportion of unseen schema vary dramatically across KBQA datasets. These discrepancies arise from the dataset construction process: WebQSP is an annotation of search logs, whereas the remaining datasets are derived from graph search and crowdsourcing. To further enhance robustness in cross-dataset transfer, we believe that better data collection methods are required to obtain diverse and balanced training data. Additionally, the representation of the logical form increases the transfer difficulty, as the S-expression used in the GrailQA dataset cannot express all queries in WebQSP.
6.4 Analysis of Modal Aspect
Context Alone Is Insufficient
We evaluate the performance of GPT-3.5 using retrieved KB contexts (§5.1.2) and in-context learning on the GrailQA dataset. The prompts for the model include the task description and the few-shot KB contexts. As illustrated in Table 7, when provided with contexts from the TIARA+GAIN retrievers, GPT-3.5 outperforms two compared models but falls short compared to TIARA+GAIN. Among the GPT-3.5 predictions, 79.62% come directly from the substring of the corresponding prompts, achieving an average F1 score of 86.19% for this portion. However, the remaining predictions are not part of their prompts and are entirely new predictions generated by GPT-3.5, with an average F1 score of merely 30.29%. Although a baseline level is attained, these results suggest that GPT-3.5 cannot be accurately grounded to the KB environment when it does not copy the retrievers’ contexts. It also shows the modal severance of natural language pre-training and KB contexts for the LLM. LLMs are not able to utilize KB contexts as proficiently as they understand natural language, and the faithfulness and controllability of grounded LLMs are not yet guaranteed under the current approach Gu et al. (2022a). To mitigate this problem, alternative paradigms should be explored, such as tool learning Schick et al. (2023) and multi-step planning Liu et al. (2023) with Chain-of-Thought Wei et al. (2022), which enables more refined access and control over environments and reduces modal differences during the reasoning process.
GrailQA GraphQ WebQSP SQB
Train size 44,337 2,381 3,097 75,819
Valid size 6,763 - - 11,141
Test size 13,231 2,395 1,638 21,483
Length 62.96 54.62 35.93 42.16
# of entities 0.903 1.028 1.112 1.000
# of relations 1.358 1.434 1.464 1.000
Similarity between questions and KB items
Entity 0.999 1.000 0.921 0.985
Class 0.547 0.457 - -
Relation 0.470 0.389 0.300 0.779
Unseen ratio (%)
Schema 16.90 86.78 20.44 32.67
Question 54.06 98.25 4.03 49.18
Table 8: KBQA dataset statistics. Length denotes the average number of question characters. # of entities/relations denotes the average number of entities/relations in the logical form. Unseen Schema is the ratio of unseen schema items in the dev/test set. Unseen Question is the ratio of questions containing unseen schema.
7 Conclusion
Despite the recent progress of LM-driven models, robustness challenges posed by distribution shifts for the KBQA task are rarely discussed. Our analyses call for further research into better evaluation protocols for grounding LMs to KBs and enhancing the robustness of environmental, linguistic, and modal aspects. Notably, the experiments reveal that LLMs sometimes simply copy the provided prompt. It indicates that the existing methodologies for grounding LLMs are yet to prove their efficacy and superiority. Future research issues include collecting more balanced environment-specific corpora and improving the LLM learning paradigms. For the corpora, our experiments show that the data augmentation techniques deserve further research.
Limitations
1) For question generation, the verbalization process of the GAIN method relies heavily on large-scale KBQA annotations. The training data influence the style of generated questions, and overly complex logical forms (e.g., with three or more hops) are difficult to convert into natural language questions. Besides, synthetic data is less diverse and natural than human annotations, though it improves generalization performance. 2) Multilingual KBQA is a problem that lacks attention in the KBQA research and is also a linguistic-aspect challenge. However, since most KBQA datasets are based on English and do not discuss other languages at all, this paper leaves the evaluation of this problem for future work.
Ethics Statement
The proposed data augmentation method GAIN could be used on any KB. The Freebase Bollacker et al. (2008) used in this work is a KB that has been publicly released and manually reviewed. For uncensored KBs, if harmful information is collected, it could make synthetic data contain harmful information and make LMs generate harmful answers.
Acknowledgments
We extend our sincere gratitude to all the anonymous reviewers for their insightful suggestions. Special thanks to Xiao Xu, Xiang Huang, and Sitao Cheng for their valuable discussions on this paper. We appreciate Yu Gu’s efforts in evaluating our submissions on the GrailQA benchmark test set.
Appendix A Example of LLM Prompt
We present an example of an excerpted prompt, which is shown in Figures 2 and 3. In particular, Figure 2 illustrates the task instruction and teaching example segment, while Figure 3 displays the input query segment.
#question #one-hop #two-hop #domain
127,329 78,668 48,661 759
#none #count #comparatives #superlatives
115,221 7,115 1,874 3,119
#class #relation #entity
5,078 12,942 46,645
Table 11: Statistics for the synthetic dataset of logical forms. none denotes no function.
#question #relation #subject #domain
162,557 7,349 108,804 673
Table 12: Statistics for the synthetic dataset of triples. Subject denotes subject entities.
Appendix B Details of Synthetic Data
The graph search process of GAIN is slightly different for logical forms and triples.
Searching Logical Forms
GAIN employs a graph search approach similar to GraphQuestions Su et al. (2016) to collect logical forms derived from graph queries. The graph query construction process consists of four steps: 1) query template construction, 2) aligning template nodes, 3) configuring functions, and 4) verification by execution. Query templates, obtained through random graph searching, contain nodes that represent entity/literal types (not aligned to a value), referred to as template nodes. Each unaligned node in the query template is then aligned with a topic entity or literal to generate multiple aligned graph queries. To synthesize counting, comparative, and superlative questions and enhance the diversity of synthetic data, we add functions like COUNT, ARGMIN/ARGMAX Gu et al. (2021) to graph queries. Since KBQA research typically assumes that questions can be answered by the KB, we execute all resulting graph queries after the above steps and remove those with a null [^11] Null for querying entities and zero for counting. result.
Searching Triples
A single KB triple can be treated as a QA pair, where the head entity and relation together form the query, and the tail entity is the answer. The triple search process consists of two steps: 1) candidate relation selection, and 2) triple sampling. First, arbitrary relations $\mathcal{R}$ are selected from the KB, without being restricted to any particular KBQA dataset. Then, triples are collected from head entities $\mathcal{H}$, where entities in $\mathcal{H}$ are connected to relations in $\mathcal{R}$.
Examples of Synthetic Data
We present some examples of synthetic data in Table 15, where the logical form contains a variety of functions.
Statistics
The statistics for KBQA datasets, including our synthetic dataset, are shown in Table 16. To calculate the number of domains in Table 11, 12, we take the domain of each Freebase class from its first segment, except for classes starting with “base”, where we take the first two segments, e.g., domain “music” for the class “music.performance_venue”, and domain “base.plants” for the class “base.plants.plant”.
Metrics GrailQA Dev GraphQ Test SQB Dev
BLEU-4 0.347 0.178 0.369
ROUGE-L 0.526 0.411 0.640
Table 13: The performance of the question generator on KBQA datasets. The generator for logical form is evaluated on the GrailQA valid set and GraphQuestions test set. The generator for triple is evaluated on the SQB valid set.
Logical Form (S-expression) Question
1. (AND book.journal (JOIN book.periodical.editorial_staff
(AND (JOIN book.editorial_tenure.editor m.012z2ncg)
(JOIN book.editorial_tenure.title m.02h6676))))
(GrailQA valid set)
2. (AND book.journal (JOIN book.periodical.editorial_staff
(AND (JOIN book.editorial_tenure.editor m.05ws_t6)
(JOIN book.editorial_tenure.title m.02wk2cy))))
(GrailQA valid set)
1. john oliver la gorce was the editor on the editor for what journal?
2. with which journal did don slater serve as editor on the editor in chief?
All four S-expressions are (COUNT (AND book.reviewed_work
(JOIN book.reviewed_work.reviews_of_this_work m.0240y2)))
(GraphQuestions training set)
1. how many works did fresh air review?
2. how many works were reviewed by fresh air in total?
3. what is the total amount of works reviewed by fresh air?
4. fresh air has reviewed how many different works?
Table 14: Examples of paraphrases in GrailQA and GraphQuestions.
Sampled Logical Form Synthetic Question
(COUNT (AND people.profession (JOIN
people.profession.people_with_this_profession m.012d40)))
|entity|m.012d40 jackie chan
how many professions does jackie chan have?
(AND food.beer (le food.beer.original_gravity 1.067^^float)) which beer has an original gravity less than or equal to 1.067?
(AND medicine.manufactured_drug_form (AND (lt
medicine.manufactured_drug_form.size 10.0^^float)
(JOIN medicine.manufactured_drug_form.fda_otc_part m.0h9yt7z)))
|entity|m.0h9yt7z fda otc monograph part 348
which manufactured drug form has part
fda otc monograph part 348 and has a size smaller than 10.0?
(ARGMAX (AND measurement_unit.power_unit
(JOIN measurement_unit.power_unit.measurement_system m.07y37))
measurement_unit.power_unit.power_in_watts)
|entity|m.07y37 us customary units
what is the largest power unit in the us customary units?
(AND music.release (AND (JOIN music.release.engineers m.011mbx12)
(JOIN music.release.label m.0g12fn3)))
|entity|m.011mbx12 raynard glass |m.0g12fn3 hostyle gospel ministries
what musical release is engineered by raynard glass and
labelled hostyle gospel ministries?
Sampled Triple Synthetic Question
D.W. Robertson, Jr. (m.09ggymq),
people.person.place_of_birth,
Washington, D.C. (m.0rh6k)
where was D. W. Robertson, Jr. born
Alfred Chao (m.046cmd8),
computer.operating_system_developer.operating_systems_developed,
pSOS (m.0lscq)
what operating system did Alfred Chao develop?
Table 15: Examples of synthetic data. The logical form is S-expression Gu et al. (2021). The entity label is appended to the logical form.
Datasets #question #class #relation #entity
GrailQA Gu et al. (2021) 64,331 1,534 3,720 32,585
GraphQuestions Su et al. (2016) 5,166 506 596 376
WebQSP Yih et al. (2016) 4,737 408 661 2,593
GAIN-synthetic 127,329 5,078 12,942 46,645
Table 16: Statistics of KBQA datasets and the GAIN-synthetic dataset.
Appendix C Experimental Details
C.1 Question Generation
Training a question generation (QG) model is the second step of GAIN. Because logical forms in the synthetic dataset do not have corresponding human-labeled natural language questions, evaluating QG on the synthetic dataset is difficult. Existing KBQA datasets contain questions written by crowdsourced workers for logical forms or triples, so we evaluate the question generator by generating questions on these datasets and calculate the BLEU-4 Papineni et al. (2002) and ROUGE-L Lin and Och (2004) scores (Table 13). [^12] Calculated by Hugging Face Evaluate.
C.2 Retrieval Performance
The performance of schema retrieval on the GrailQA valid set and GraphQuestions test set is shown in Table 9 and 10, respectively. GAIN improves the performance of schema retrieval on both GrailQA and GraphQuestions. In particular, GAIN improves the relation recall@10 by 7.2 points on GraphQuestions.
Function None Count Comparative Superlative
ArcaneQA Gu and Su (2022) 70.8/77.8 62.5/68.2 54.5/75.7 70.5/75.6
RnG-KBQA Ye et al. (2022) 77.5/81.8 73.0/77.5 55.1/76.0 13.8/22.3
TIARA (T5-base) Shu et al. (2022) 77.8/83.1 76.4/81.8 57.4/81.4 58.7/69.0
TIARA + GAIN ELF only 76.8/81.7 73.9/80.0 0.0/25.3 0.0/8.3
TIARA + GAIN (T5-base) 78.6/84.6 77.7/83.0 61.7/82.3 69.9/73.2
TIARA + GAIN (T5-3B) 78.5/84.8 77.3/82.5 63.0/84.5 70.7/74.1
GPT-3.5-turbo (5-shot) 74.1/78.0 66.8/70.5 38.3/60.5 43.9/52.3
# of relations 1 2 3 4
RnG-KBQA Ye et al. (2022) 75.7/79.3 65.3/74.7 28.6/44.5 100.0/100.0
TIARA (T5-base) Shu et al. (2022) 81.2/85.6 64.7/75.8 29.3/48.5 50.0/83.3
TIARA + GAIN ELF only 74.0/77.8 56.6/67.9 9.9/31.0 0.0/33.3
TIARA + GAIN (T5-base) 82.4/87.2 67.0/78.0 38.9/49.8 50.0/83.3
TIARA + GAIN (T5-3B) 82.0/87.2 68.8/79.0 37.5/51.3 50.0/83.3
GPT-3.5-turbo (5-shot) 75.0/78.6 61.9/69.6 19.8/36.6 50.0/50.0
# of entities 0 1 2
RnG-KBQA Ye et al. (2022) 58.5/63.6 75.4/79.9 55.6/73.5
TIARA (T5-base) Shu et al. (2022) 77.5/83.1 76.6/82.6 49.9/68.0
TIARA + GAIN ELF only 42.8/47.0 74.2/79.9 47.4/67.6
TIARA + GAIN (T5-base) 82.2/86.5 77.6/83.7 53.4/71.4
TIARA + GAIN (T5-3B) 82.0/86.6 77.6/84.0 55.6/73.0
GPT-3.5-turbo (5-shot) 66.8/72.8 71.9/76.1 48.8/62.1
Table 17: EM and F1 scores (%) for different types of logical forms on the GrailQA valid set. None denotes no function. # of relations/entities denotes the number of relations/entities in the S-expression. ELF denotes exemplary logical form Shu et al. (2022).
C.3 Performance on Various Logical Forms
To show how TIARA+GAIN performs on different types of logical forms, we compare it with previous KBQA models on the GrailQA valid set, as shown in Table 17. TIARA+GAIN improves performance in nearly all these scenarios compared to TIARA.
C.4 Error Analysis
To analyze the QA errors of TIARA+GAIN (T5-3B), we randomly sample 50 questions where predicted logical forms are not the same as the ground truth in the GrailQA valid set. We follow Shu et al. (2022) in their classification of errors. Entity linking errors (missing or redundant entities), syntactic errors (generating the wrong logical form structure), semantic errors (generating the wrong schema item when correct contexts are provided), false negatives (flawed annotations), and miscellaneous (e.g., ambiguity in the question) account for 48%, 26%, 16%, 4%, and 6%, respectively. For entity linking errors, 62.5% of them are from the zero-shot level. For syntactic errors, the number of errors from zero-shot and compositional levels is almost the same. It means that entity linking for zero-shot domains and logical form generation for complex structures remain significant challenges.
Appendix D Details of Scientific Artifacts
All datasets we use are publicly available. GrailQA [^13] https://dki-lab.github.io/GrailQA Gu et al. (2021) uses CC BY-SA 4.0 license, and GraphQuestions [^14] https://github.com/ysu1989/GraphQuestions Su et al. (2016) uses CC BY 4.0 license. WebQSP [^15] https://www.microsoft.com/en-us/download/details.aspx?id=52763 Yih et al. (2016) and SimpleQuestions-Balance [^16] https://github.com/wudapeng268/KBQA-Adapter/tree/master/Data/SQB Wu et al. (2019) are also downloaded from their official release channels. We have complied with their distribution rules. These datasets involve manual construction rather than fully automated construction, which includes the review process. They contain questions about some famous individual people, but the corresponding content is available on Freebase, which is a publicly released, human-reviewed knowledge base.
Although the training data for LLMs could be quite large, the test set annotation of GrailQA is not publicly available. In addition, our experimental results on the test set and the validation set show the same trend, so the impact of the data contamination problem on the experiments of this paper could be ignored.
Given a question and Freebase contexts, write a logical form that answers the question. Question: dark sun: wake of the ravager was designed by what video game designer? Candidate entities: (A) [Dark Sun: Wake of the Ravager] Exemplary Logical Forms: (A) (AND cvg.cvg_designer (JOIN cvg.cvg_designer.games_designed [Dark Sun: Wake of the Ravager])) (B) (AND cvg.cvg_designer (JOIN (R cvg.computer_videogame.designers) [Dark Sun: Wake of the Ravager])) (C) (AND cvg.computer_videogame (JOIN (R cvg.cvg_designer.games_designed) (JOIN cvg.cvg_designer.games_designed [Dark Sun: Wake of the Ravager]))) (D) (AND cvg.computer_videogame (JOIN cvg.computer_videogame.designers (JOIN cvg.cvg_designer.games_designed [Dark Sun: Wake of the Ravager]))) (E) (AND base.wikipedia_infobox.video_game (JOIN base.wikipedia_infobox.video_game.developer (JOIN cvg.cvg_designer.games_designed [Dark Sun: Wake of the Ravager]))) Candidate classes: (A) cvg.cvg_designer (B) cvg.game_performance (C) cvg.musical_game (D) cvg.game_character (E) cvg.computer_game_engine_developer (F) cvg.computer_videogame (G) cvg.computer_game_performance_type (H) cvg.game_version (I) cvg.computer_game_subject (J) cvg.computer_game_evaluation Candidate relations: (A) cvg.cvg_designer.games_designed (B) cvg.computer_videogame.designers (C) cvg.computer_videogame.prequel (D) cvg.computer_videogame.sequel (E) cvg.computer_videogame.mods (F) cvg.computer_videogame.expansions (G) cvg.computer_videogame.developer (H) cvg.computer_videogame.characters (I) cvg.game_version.game (J) cvg.computer_game_mod.game_modded Prediction: (AND cvg.cvg_designer (JOIN cvg.cvg_designer.games_designed [Dark Sun: Wake of the Ravager]))
Figure 2: Example of LLM prompt (part 1): The task instruction and $k$ teaching examples (only one is shown because of the length) with questions and their contexts.
Question: worldofwarcraft is the creation of which video game designer? Candidate entities: (A) [worldofwarcraft] Exemplary Logical Forms: (A) (AND cvg.cvg_designer (JOIN cvg.cvg_designer.games_designed [worldofwarcraft])) (B) (AND cvg.cvg_designer (JOIN (R cvg.computer_videogame.designers) [worldofwarcraft])) (C) (AND cvg.cvg_designer (JOIN cvg.cvg_designer.games_designed (JOIN cvg.computer_game_expansion.expansion_for [worldofwarcraft]))) (D) (AND cvg.cvg_designer (JOIN (R cvg.computer_videogame.designers) (JOIN cvg.computer_game_expansion.expansion_for [worldofwarcraft]))) (E) (AND cvg.cvg_designer (JOIN (R cvg.computer_videogame.designers) (JOIN (R cvg.computer_videogame.expansions) [worldofwarcraft]))) Candidate classes: (A) games.game_designer (B) cvg.cvg_designer (C) amusement_parks.ride_designer (D) cvg.cvg_developer (E) cvg.computer_videogame (F) cvg.computer_game_engine_developer (G) cvg.computer_game_engine (H) cvg.computer_game_mod (I) cvg.game_performance (J) cvg.musical_game Candidate relations: (A) cvg.computer_videogame.designers (B) cvg.cvg_designer.games_designed (C) games.game_designer.games_designed (D) games.game.designer (E) cvg.computer_videogame.developer (F) cvg.cvg_developer.games_developed (G) cvg.computer_game_engine.developer (H) cvg.computer_videogame.expansions (I) cvg.computer_videogame.publisher (J) cvg.game_version.developer Prediction:
Figure 3: Example of LLM prompt (part 2): After $k$ teaching examples, the input question and its contexts is given. The golden prediction is (AND cvg.cvg_designer (JOIN cvg.cvg_designer.games_designed [worldofwarcraft])).